Garcia Industries has sales of $200,000 and accounts receivable of $18,500, and it gives its customers 25 days to pay. The industry average DSO is 27 days, based on a 365-day year. If the company changes its credit and collection policy sufficiently to cause its DSO to fall to the industry average, and if it earns 8.0% on any cash freed-up by this change, how would that affect its net income, assuming other things are held constant?
Answer #1

Present DSO = (Accounts receivable/ Total credit sales) * number of days

= ($18,500 / $200,000) * 365 = 0.0925 * 365 = 33.76 days

Company wants to reduce its DSO to Industry Average of 27 Days

DSO = (Accounts receivable/Total credit sales) * number of days

27= (Expected accounts receivable / $200,000) * 365

Expected accounts receivable= (27* $200,000)/365 = $14,794.52

Therefore the decrease in accounts receivable due to the change in policy would be

= $18,500 - $14,794.52 = $3,705.48

This policy would free up cash of $3,705.48, on which the company will earn 8%.

The income earned = $3,705.48 * 8% = $296.44

So, the net income would increase by $296.44
Answer #2
The days sales outstanding (DSO) formula is:
(Accounts Receivable / Credit Sales) x 365 = DSO

Currently, Aziz Industries DSO is:
(18,500 / 200,000) x 365 = 33.76 Days

The question is asking you, what is the amount outstanding Accounts Receivable needs to be lowered to in order for DSO to be 27 days. And how much interest you would earn on the freed up cash. To calculate what Accounts Receivable needs to be lowered to, use the DSO formula and insert x for Accounts Receivable and use 27 as the number of days.

(x / 200,000) x 365 = 27
Solve for x
x = $14,794.52 This is the amount Accounts Receivable needs to be lowered to.
18,500 - 14,794.52 = 3,705.48 decrease in Accounts Receivable
3,705.48 x 8% = $296.43 Interest eaned and addition to net income.

Answer is $296.43

PLEASE RATE IF IT HELPS YOU :)
Answer #3
Rate of return on cash generated 8.0%

Sales $200,000

A/R $18,500

Days in Year 365

Sales/day = Sales/365 = $547.95

Company DSO = Receivables/Sales per day = 33.8

Industry DSO 27.0

Difference = Company DSO – Industry DSO = 6.8

Cash flow from reducing the DSO = Difference × Sales/day = $3,705.48

Additional Net Income = Return on cash × Added cash flow = $296.44



Alternative Calculation:

A/R at industry DSO $14,794.52

Change in A/R $3,705.48

Additional Net Income $296.44